# Detect, Investigate, Judge and Determine: A Knowledge-guided Framework for Few-shot Fake News Detection

### v3

@@ Abstract
Detect, Investigate, Judge and Determine:
A Knowledge-guided Framework for Few-shot Fake News Detection
Ye Liu1,2, Jiajun Zhu1,2, Xukai Liu1,2, Haoyu Tang1,2 Yanghai Zhang1,2, Kai Zhang1,2, Xiaofang Zhou3, Enhong Chen1,2 1 University of Science and Technology of China 2 State Key Laboratory of Cognitive Intelligence 3 The Hong Kong University of Science and Technology liuyer@mail.ustc.edu.cn
1 Introduction
Fake News Detection (FND), aiming to distinguish between inaccurate news and legitimate news, has garnered increasing importance and attention due to the the pervasive dissemination and detrimental effects of fake news on social media platforms (Shu et al., 2017). Few-Shot Fake News Detection (FS-FND), as a subtask of FND, endeavors to identify fake news by leveraging only $K$ instances per category ($K$-shot) in the training phase (Hu et al., 2024; Gao et al., 2021; Ma et al., 2023).
@@ 1 Introduction
Generally, fake news detection can be framed as a binary classification problem and addressed using various classification models. In the early stage, researchers primarily employ machine learning or deep learning algorithms to represent and classify candidate news articles (Horne and Adali, 2017; Jiang et al., 2022). More recently, with the rise of Large Language Models (LLMs), FSFND has been effectively addressed through the in-context learning technology, which is particularly prevalent in few-shot settings (Hu et al., 2024; Boissonneault and Hensen, 2024). Among them, Hu et al. (2024); Wang et al. (2024); Teo et al. (2024) were pioneers in investigating the potential of LLMs in this field.
@@
2 Related Work
2.1 Few-Shot Fake News Detection
@@
Generally, fake news detection can be defined as a binary classification problem and addressed by a variety of classification models. Initially, researchers mainly rely on feature engineering and machine learning algorithms (Horne and Adali, 2017). As computing power and data availability have increased, significant improvements have been made with the help of various deep learning algorithms and Pre-trained Language Models (PLMs). For instance, Ghanem et al. (2021) combined lexical features and a Bi-GRU network to achieve accurate fake news detection. Jiang et al. (2022) introduced the Knowledgeable Prompt Learning (KPL), a novel framework that integrates prompt learning with fake news detection, achieving state-of-the-art performance in few-shot settings.
@@
Additionally, researchers have also recognized the importance of external knowledge to complement traditional fake news detection methods. News articles often contain references to entities, events, and facts that are external to the article itself, which makes external knowledge integration crucial for accurate detection. For instance, Dun et al. (2021); Ma et al. (2023); Hu et al. (2021b) utilized knowledge graphs to enrich entity information and structured relation knowledge, leading to more precise news representations and improved detection performance. Meanwhile, Huang et al. (2023) adopted a data augmentation perspective, proposing a novel framework for generating more valuable training examples, which has proven to be beneficial in detecting human-written fake news.
@@
More recently, with the advent of large language models, many researchers are exploring few-shot fake news detection through in-context learning and data augmentation technologies (Hu et al., 2024; Wang et al., 2024; Teo et al., 2024). For example, Hu et al. (2024) explored the role of LLMs in fake news detection, proposing an Adaptive Rationale Guidance (ARG) network that combines traditional detection techniques with the generative capabilities of LLMs. However, many of these approaches rely on directly prompting the LLM to classify the news as fake or real, without fully utilizing the broader potential of LLMs to understand and reason about the news context. More importantly, most of them are significantly limited by the aforementioned two shortcomings, particularly in the Information Scarcity problem.
2.2 Large Language Models & Retrieval-Augmented Generation
@@
Large Language Models (LLMs) such as GPT-4, LLama-3, and others have revolutionized natural language processing, achieving impressive results across a variety of tasks, including text classification, summarization, information extraction, and fake news detection (Hoffmann et al., 2022; OpenAI, 2023; AI@Meta, 2024; Tunstall et al., 2023). One of the key innovations introduced by LLMs is in-context learning, a paradigm in which the model learns from a few examples presented in the context of the task, allowing it to perform few-shot learning without the need for explicit retraining (Brown et al., 2020; Hu et al., 2024; Wang et al., 2024).
@@
In parallel, the integration of Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance the capabilities of LLMs. RAG models, such as those proposed by Lewis et al. (2020), combine the generative power of LLMs with the retrieval of relevant documents or knowledge from an external corpus. This approach not only improves the factual accuracy of generated content but also allows for more informed and contextually appropriate responses in tasks like question answering and knowledge-intensive applications (Izacard and Grave, 2021; Borgeaud et al., 2022; Izacard et al., 2023).
@@
In the context of few-shot fake news detection, incorporating external knowledge retrieval has shown promising results. Jiang et al. (2022) were among the first to integrate retrieval-based methods with LLMs for fake news detection. By incorporating knowledge from Wikidata into the prompt representation process, they devised a knowledgeable prompt learning framework that significantly improved the detection performance. Similarly, other studies have optimized external retrieval mechanisms to enhance the performance of LLMs, allowing more context-aware and accurate predictions (Borgeaud et al., 2022; Izacard et al., 2023).
3 Problem Statement
@@
Generally, fake news detection can be framed as a binary classification problem, wherein each news article is classified as either real ($y=0$) or fake ($y=1$(Dun et al., 2021). Formally, each piece of news $S$ is composed of a sequence of words, i.e., $S=\{s_{1},s_{2},...,s_{n}\}$, encompassing its title, content text, and relevant tweets. The goal is to learn a detection function $F:F(S)\Longrightarrow y$, where $y\in\{0,1\}$ denotes the ground-truth label of news.
@@
In the few-shot settings, adhering the strategy employed in (Gao et al., 2021; Ma et al., 2023), we randomly sample $K$ instances per category ($K$-shot) [^1] This implies that for a $K$-shot fake news detection setting, the number of training instances is $2K$. for the training phase. The entire test set is preserved to ensure the comprehensiveness and effectiveness of evaluation.
4 The DKFND Model
@@
4.1 Detection Module
In this module, we aim to identify the key information contained in the given news article, which will serve as the query for the subsequent modules. Specifically, we employ the Spacy algorithm Honnibal and Montani (2017) to extract relevant knowledge concepts $C=\{c_{1},c_{2},...,c_{m}\}$.
@@ 4.1 Detection Module
where $\mathcal{F}$ represents the LLM, and $P_{detection}$ denotes the prompt for the in-context learning. More detailed description can be found in Appendix C.1.
4.2 Investigation Module
@@
4.2.1 Inside Investigation.
To address the Understanding Ambiguity problem introduced in Section 1, we retrieve effective demonstrations to enhance the LLM’s understanding during the in-context learning process (Liu et al., 2022; Rubin et al., 2022).
@@
Subsequently, when inferring a candidate news $j$, we employ the $k$-Nearest Neighbors ($k$NN) search method (Khandelwal et al., 2019) to retrieve valuable samples from the training set. In detail, we use the representation $H_{j}$ of news $j$ to query the datastore $D$ according to the euclidean distance. Then, based on the computed distance, we select the nearest $k$ positive and negative news samples, respectively:
@@
4.2.2 Outside Investigation.
In response to the Information Scarcity problem, we further retrieve additional real-time information from external sources. Inspired by Yoran et al. (2023); Paranjape et al. (2023), we implement a retriever based on the google search engine, using the SerpAPI service [^2] https://serpapi.com/. You can refer to Section B for a searching example..
Specifically, based on the extracted concetps $\{c_{1},c_{2},...,c_{N}\}$ in Section 4.1, we first concatenate them to construct the initial query $\mathcal{Q}=c_{1}\uplus c_{2}\uplus...\uplus c_{N}$. Then, following the strategy proposed in (Yoran et al., 2023), we further format the search queries as “en.wikipedia.org $\mathcal{Q}$”, with the Wikipedia domain preceding the intermediate question. We return the top-L evidence retrieved by Google. And all retrieved evidence sentences are prepended to the outside investigation outcome, denoted as $\mathcal{OI}=\{G_{k},k=1,2,..,L\}$.
4.3 Judge Module
@@
4.3.1 Inside Judge
In the Inside Investigation module (Section 4.2.1), we employ the kNN search method to retrieve relevant documents $\mathcal{II}$ from the training set. However, the retrieved content may still contain irrelevant information, potentially introducing noise and interfering with the prediction process.
@@ 4.3.1 Inside Judge
where $P_{IJ}$ denotes the prompt and more detailed can be found in Appendix C.2. Finally, the selected inside investigated information can be denoted as =$\mathcal{II}^{\prime}=\{U_{positive}^{\prime},U_{negative}^{\prime}\}$.
4.3.2 Outside Judge
For the Ouside Investigation module (Section 4.2.2), although the retrieved information is obtained via the Google API, ensuring the relevance, a major limitation lies in the uncertainty of its authenticity and accuracy. To address this, we incorporate a knowledge graph to verify the authenticity of the retrieved outside information.
@@ 4.3.2 Outside Judge
In this part, we further employ the KG embedding model: ComplEx (Trouillon et al., 2016), which is pretrained on Wikidata through LibKGE (Broscheit et al., 2020), to assess the authenticity of each triple $T_{k}^{l}$ in $\mathcal{T}_{k}$:
@@
Relying on the authenticity score of $G_{k}\in\mathcal{OI}$, we could finally obtain the selected documents. Specifically, considering the ranking design of Google search (Patil et al., 2021), we first choose the top-1 document $G_{1}$ in $\mathcal{OI}$. Then we select top n-1 documents from the remaining L-1 documents, through the authenticity score $Score(G_{k})$. Finally, we obtain the selected outside investigated information $\mathcal{OI^{\prime}}=\{G_{k},k=1,2,..,n\}$.
4.4 Determination Module
4.4.1 Inside Determination
@@
where $P_{ID}$ is the prompt instruction, $P_{i}$ refers to the prediction result, while $R_{i}$ denotes the corresponding explanation. You can move to Appendix C.2 for more details about this prompt.
4.4.2 Outside Determination
@@ 4.4.3 Integrated Determination
5 Experiments
5.1 Experiment Setup
5.1.1 Datasets and Evaluation Metrics.
We conduct experiments on two datasets, PolitiFact and Gossipcop, both of which are proposed in a benchmark called FakeNewsNet (Shu et al., 2020). PolitiFact consists of various political news, while Gossipcop is sourced from an entertainment story fact-checking website. For the few-shot setting, following the strategy employed in (Jiang et al., 2022; Ma et al., 2023), we randomly select $K\in(8,32,100)$ positive and negative news articles as the training set, respectively. More statistics about the datasets are illustrated in Table 1.
Given that the task focuses on detecting fake news, fake news articles are regarded as positive examples (Ma et al., 2023). We further adopt the F1-score and Accuracy (ACC) as the evaluation metrics to measure classification performance.
@@ 5.1 Experiment Setup
Table 1: Statistics of PolitiFact and Gossipcop datasets.
@@
Table 2: Experimental results of our proposed method on the PolitiFact and Gossipcop datasets.
5.1.2 Implementation Details.
In DKFND architecture, we utilize the zephyr-7b-beta (Tunstall et al., 2023) model on Huggingface as the LLM. When running Zephyr, we adhere to the default parameter values provided by the official, where the sampling temperature is $0.70$, top_k is $50$, and top_p is $0.95$. The max_new_token is set to $256$, and do_sample is set as $True$.
In the Detection Module (Section 4.1), the number of keywords to extract is set to $N=5$.
In the Inside Investigation part (Section 4.2.1), we employ the DeBERTa-base model (He et al., 2021) from Transformers (Wolf et al., 2020) as the representation model. The number of retrieved positive/negative nearest neighbors is set as $k=5$.
In the Outside Investigation part (Section 4.2.2), we set the number of retrieved documents as $L=8$.
In the Inside Judge part (Section 4.3.1), the number of positive/negative samples is set as $m=2$.
In the Outside Judge part (Section 4.3.2), we set the number of selected documents as $n=2$.
@@
We compare DKFND with the state-of-the-art few-shot fake news detection methods. According to the model architecture, they can be grouped into three categories, including traditional fake news detection methods (① $\sim$ ⑤), LLM-based methods (⑥ $\sim$ 10), and hybrid methods ( 11).
@@
  • PROPANEWS (Huang et al., 2023) proposes a framework for generating valuable training examples, beneficial to human-written situations.
@@
  • FakeFlow (Ghanem et al., 2021) devises a model that detects fake news articles by integrating the flow of affective information.
@@
  • MDFEND (Nan et al., 2021) incorporates the domain information through a gate mechanism to aggregate multiple representations.
@@
  • PSM (Ni et al., 2020) utilizes Propensity Score Matching to select decounfounded features, boosting detection performance.
@@
  • KPL (Jiang et al., 2022) incorporates external knowledge into the prompt representation process, achieving knowledgeable detection.
@@
  • AutoCoT (Zhang et al., 2022) proposes an automatic chain-of-thought prompting method to construct demonstrations and reasoning chains.
@@
  • Zephyr (Tunstall et al., 2023) represents the advanced 7B model, which is optimized by the preference data from AI Feedback.
  • ChatGLM-3 (Zeng et al., 2022; Du et al., 2022) is a series of pre-trained dialogue models, and we select the ChatGLM3-6B version.
@@
  • LLama-3 (AI@Meta, 2024) refers to the LLM proposed by Meta. We adopt 8B version (Meta-Llama-3-8B-Instruct) for experiments.
@@
  • 10 GPT-3.5 (Ouyang et al., 2022) is an advanced LLM developed by OpenAI. We leverage the API (version: gpt-3.5-turbo-0613) for ICL.
@@
  • 11 ARG (Hu et al., 2024) designs an adaptive rationale guidance network, integrating insights from both LLMs and traditional methods.
@@
It is worth noting that, for these LLM-based baselines (⑦ $\sim$ 10), we adhere to the instruction prompt proposed by (Hu et al., 2024) to conduct in-context learning. Besides, due to the limitations of maximum tokens, we randomly select 6 samples as the demonstrations, which is more than the demonstration samples utilized in our DKFND model [^3] As introduced in Section 5.1.2, DKFND utilizes $n=2$ positive and negative samples as demonstrations, with a total number of 4., facilitating a fair comparison.
5.2 Experimental Result

### v5

@@ Abstract
Index Terms:
fake news detection, few shot, large language models
I Introduction
Fake News Detection (FND), aiming to distinguish between inaccurate news and legitimate news, has garnered increasing importance and attention due to the the pervasive dissemination and detrimental effects of fake news on social media platforms[1]. Few-Shot Fake News Detection (FS-FND), as a subtask of FND, endeavors to identify the fake news by leveraging only $K$ instances per category ($K$-shot) in the training phase[2, 3, 4].
Generally, fake news detection can be framed as a binary classification problem and addressed using various classification models. In the early stage, researchers primarily employ machine learning or deep learning algorithms to represent and classify candidate news articles[5, 6]. More recently, with the rise of Large Language Models (LLMs), FSFND has been effectively addressed through the in-context learning technology, which is particularly prevalent in few-shot settings[2, 7]. Among them, Hu and Wang et al.[2, 8, 9] were pioneers in investigating the potential of LLMs in this field.
@@ I Introduction
II Related Work
II-A Few-Shot Fake News Detection
@@
Generally, fake news detection can be defined as a binary classification problem and addressed by a variety of classification models. Initially, researchers mainly rely on feature engineering and machine learning algorithms[5]. As computing power and data availability have increased, significant improvements have been made with the help of various deep learning algorithms and Pre-trained Language Models (PLMs). For instance, Ghanem et al.[10] combined lexical features and a Bi-GRU network to achieve accurate fake news detection. Jiang et al.[6] introduced the Knowledgeable Prompt Learning (KPL), a novel framework that integrates prompt learning with fake news detection, achieving state-of-the-art performance in few-shot settings.
@@
Additionally, researchers have also recognized the importance of external knowledge to complement traditional fake news detection methods. News articles often contain references to entities, events, and facts that are external to the article itself, which makes external knowledge integration crucial for accurate detection. For instance, Dun and Ma et al.[11, 4, 12] utilized knowledge graphs to enrich entity information and structured relation knowledge, leading to more precise news representations and improved detection performance. Meanwhile, Huang et al.[13] adopted a data augmentation perspective, proposing a novel framework for generating more valuable training examples, which has proven to be beneficial in detecting human-written fake news.
@@
More recently, with the advent of large language models, many researchers are exploring few-shot fake news detection through in-context learning and data augmentation technologies[2, 8, 9]. For example, Hu et al.[2] explored the role of LLMs in fake news detection, proposing an Adaptive Rationale Guidance (ARG) network that combines traditional detection techniques with the generative capabilities of LLMs. However, many of these approaches rely on directly prompting the LLM to classify the news as fake or real, without fully utilizing the broader potential of LLMs to understand and reason about the news context. More importantly, most of them are significantly limited by the aforementioned two shortcomings, particularly in the Information Scarcity problem.
II-B Large Language Models & Retrieval-Augmented Generation (RAG)
@@
Large Language Models (LLMs) such as GPT-4, LLama-3, and others have revolutionized natural language processing, achieving impressive results across a variety of tasks, including text classification, summarization, information extraction, and fake news detection[14, 15, 16, 17]. One of the key innovations introduced by LLMs is in-context learning, a paradigm in which the model learns from a few examples presented in the context of the task, allowing it to perform few-shot learning without the need for explicit retraining[18, 2, 8].
@@
In parallel, the integration of Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance the capabilities of LLMs. RAG models, such as those proposed by[19], combine the generative power of LLMs with the retrieval of relevant documents or knowledge from an external corpus. This approach not only improves the factual accuracy of generated content but also allows for more informed and contextually appropriate responses in tasks like question answering and knowledge-intensive applications[20, 21, 22].
@@
In the context of few-shot fake news detection, incorporating external knowledge retrieval has shown promising results. Jiang et al.[6] were among the first to integrate retrieval-based methods with LLMs for fake news detection. More specifically, by incorporating knowledge from Wikidata into the prompt representation process, they devised a knowledgeable prompt learning framework that significantly improved the detection performance. Similarly, other studies have optimized external retrieval mechanisms to enhance the performance of LLMs, allowing more context-aware and accurate prediction results[21, 22].
III Problem Statement
@@
Generally, fake news detection can be framed as a binary classification problem, wherein each news article is classified as either real ($y=0$) or fake ($y=1$)[11]. Formally, each piece of news $S$ is composed of a sequence of words, i.e., $S=\{s_{1},s_{2},...,s_{n}\}$, encompassing its title, content text, and relevant tweets. The goal is to learn a detection function $F:F(S)\Longrightarrow y$, where $y\in\{0,1\}$ denotes the ground-truth label of news.
@@
In the few-shot settings, adhering the strategy employed in[3, 4], we randomly sample $K$ instances per category ($K$-shot) [^1] This implies that for a $K$-shot fake news detection setting, the number of training instances is $2K$. for the training phase. The entire test set is preserved to ensure the comprehensiveness and effectiveness of evaluation.
IV The DKFND Model
@@
IV-A Detection Module
In this module, we aim to identify the key information contained in the given news article, which will serve as the query for the subsequent modules. Specifically, we employ the Spacy algorithm[23] to extract relevant knowledge concepts $C=\{c_{1},c_{2},...,c_{m}\}$.
@@ IV-A Detection Module
where $\mathcal{F}$ represents the LLM, and $P_{detection}$ denotes the prompt for the in-context learning. More detailed description can be found in Appendix -B.
IV-B Investigation Module
@@
IV-B1 Inside Investigation.
To address the Understanding Ambiguity problem introduced in Section I, we retrieve effective demonstrations to enhance the LLM’s understanding during the in-context learning process[24, 25].
@@
Subsequently, when inferring a candidate news $j$, we employ the $k$-Nearest Neighbors ($k$NN) search method[26] to retrieve valuable samples from the training set. In detail, we use the representation $H_{j}$ of news $j$ to query the datastore $D$ according to the euclidean distance. Then, based on the computed distance, we select the nearest $k$ positive and negative news samples, respectively:
@@
IV-B2 Outside Investigation.
In response to the Information Scarcity problem, we further retrieve additional real-time information from external sources. Inspired by[27, 28], we implement a retriever based on the google search engine, using the SerpAPI service [^2] https://serpapi.com/. You can refer to Appendix -A for a searching example..
Specifically, based on the extracted concetps $\{c_{1},c_{2},...,c_{N}\}$ in Section IV-A, we first concatenate them to construct the initial query $\mathcal{Q}=c_{1}\uplus c_{2}\uplus...\uplus c_{N}$. Then, following the strategy proposed in[27], we further format the search queries as “en.wikipedia.org $\mathcal{Q}$”, with the Wikipedia domain preceding the intermediate question. We return the top-L evidence retrieved by Google. And all retrieved evidence sentences are prepended to the outside investigation outcome, denoted as $\mathcal{OI}=\{G_{k},k=1,2,..,L\}$.
IV-C Judge Module
@@
IV-C1 Inside Judge
In the Inside Investigation module (Section IV-B1), we employ the kNN search method to retrieve relevant documents $\mathcal{II}$ from the training set. However, the retrieved content may still contain irrelevant information, potentially introducing noise and interfering with the prediction process.
@@ IV-C1 Inside Judge
where $P_{IJ}$ denotes the prompt and more detailed can be found in Appendix -B. Finally, the selected inside investigated information can be denoted as =$\mathcal{II}^{\prime}=\{U_{positive}^{\prime},U_{negative}^{\prime}\}$.
IV-C2 Outside Judge
For the Ouside Investigation module (Section IV-B2), although the retrieved information is obtained via the Google API, ensuring the relevance, a major limitation lies in the uncertainty of its authenticity and accuracy. To address this, we incorporate a knowledge graph to verify the authenticity of the retrieved outside information.
@@ IV-C2 Outside Judge
In this part, we further employ the KG embedding model: ComplEx[29], which is pretrained on Wikidata through LibKGE[30], to assess the authenticity of each triple $T_{k}^{l}$ in $\mathcal{T}_{k}$:
@@
Relying on the authenticity score of $G_{k}\in\mathcal{OI}$, we could finally obtain the selected documents. Specifically, considering the ranking design of Google search[31], we first choose the top-1 document $G_{1}$ in $\mathcal{OI}$. Then we select top n-1 documents from the remaining L-1 documents, through the authenticity score $Score(G_{k})$. Finally, we obtain the selected outside investigated information $\mathcal{OI^{\prime}}=\{G_{k},k=1,2,..,n\}$.
IV-D Determination Module
IV-D1 Inside Determination
@@
where $P_{ID}$ is the prompt instruction, $P_{i}$ refers to the prediction result, while $R_{i}$ denotes the corresponding explanation. You can move to Appendix -B for more details about this prompt.
IV-D2 Outside Determination
@@ IV-D3 Integrated Determination
V Experiments
V-A Experiment Setup
V-A1 Datasets and Evaluation Metrics.
We conduct experiments on two datasets, PolitiFact and Gossipcop, both of which are proposed in a benchmark called FakeNewsNet[32]. PolitiFact consists of various political news, while Gossipcop is sourced from an entertainment story fact-checking website. For the few-shot setting, following the strategy employed in[6, 4], we randomly select $K\in(8,32,100)$ positive and negative news articles as the training set, respectively. More statistics about the datasets are illustrated in Table I.
Given that the task focuses on detecting fake news, fake news articles are regarded as positive examples[4]. We further adopt the F1-score and Accuracy (ACC) as the evaluation metrics to measure classification performance.
TABLE I: Statistics of PolitiFact and Gossipcop datasets.
@@ V-A Experiment Setup
TABLE II: Experimental results of our proposed method on the PolitiFact and Gossipcop datasets.
@@
V-A2 Implementation Details.
In DKFND architecture, we utilize the zephyr-7b-beta[17] model on Huggingface as the LLM. When running Zephyr, we adhere to the default parameter values provided by the official, where the sampling temperature is $0.70$, top_k is $50$, and top_p is $0.95$. The max_new_token is set to $256$, and do_sample is set as $True$.
In the Detection Module (Section IV-A), the number of keywords to extract is set to $N=5$.
In the Inside Investigation part (Section IV-B1), we employ the DeBERTa-base model[33] from Transformers[34] as the representation model. The number of retrieved positive/negative nearest neighbors is set as $k=5$.
In the Outside Investigation part (Section IV-B2), we set the number of retrieved documents as $L=8$.
In the Inside Judge part (Section IV-C1), the number of positive/negative samples is set as $m=2$.
In the Outside Judge part (Section IV-C2), we set the number of selected documents as $n=2$.
@@
For demonstrating DKFND’s effectiveness, we compare it with the state-of-the-art few-shot fake news detection methods. According to the model architecture, they can be grouped into three categories, including traditional fake news detection methods (① $\sim$ ⑤), LLM-based methods (⑥ $\sim$ 10), and hybrid methods ( 11).
@@
  1. PROPANEWS[13] proposes a framework for generating valuable training examples, beneficial to human-written situations.
@@
  2. FakeFlow[10] devises a model that detects fake news articles by integrating the flow of affective information.
@@
  3. MDFEND[35] incorporates the domain information through a gate mechanism to aggregate multiple representations.
@@
  4. PSM[36] utilizes Propensity Score Matching to select decounfounded features, boosting detection performance.
@@
  5. KPL[6] incorporates external knowledge into the prompt representation process, thus achieving knowledgeable detection for fake news.
@@
  6. AutoCoT[37] proposes an automatic chain-of-thought prompting method to construct demonstrations and reasoning chains.
@@
  7. Zephyr[17] represents the advanced 7B model, which is optimized by the preference data from AI Feedback.
  8. ChatGLM-3[38, 39] is a series of pre-trained dialogue models, and we select the ChatGLM3-6B version.
@@
  9. LLama-3[16] refers to the LLM proposed by Meta AI. We adopt its 8B version (Meta-Llama-3-8B-Instruct) for experiments.
  10. 10
@@
    GPT-3.5[40] is an advanced LLM developed by OpenAI. We leverage the API (version: gpt-3.5-turbo-0613) for in-context learning.
  11. 11
@@
    ARG[2] designs an adaptive rationale guidance network, which integrates insights from both large language models and traditional methods.
@@
It is worth noting that, for these LLM-based baselines (⑦ $\sim$ 10), we adhere to the instruction prompt proposed by[2] to conduct in-context learning. Besides, due to the limitations of maximum tokens, we randomly select 6 samples as the demonstrations, which is more than the demonstration samples utilized in our DKFND model [^3] As introduced in Section V-A2, DKFND utilizes $n=2$ positive and negative samples as demonstrations, with a total number of 4., facilitating a fair comparison.
V-B Experimental Result